# Chicago, IL, USA

### v1

@@ 1 Introduction
As manually finding vulnerabilities typically incur considerable effort and time, numerous studies have been dedicated to automatically identify vulnerabilities (Kim et al., 2017; Li et al., 2016; Grieco et al., 2016; Neuhaus et al., 2007a; Yamaguchi et al., 2013, 2012a). Primarily, these approaches rely on code similarity detection or pattern matching techniques. However, code similarity detection may not effectively identify vulnerabilities that do not result from code duplication, and pattern matching necessitates the expertise of human professionals to define vulnerability patterns.
@@ 5 Related Work
Many techniques have been proposed to detect and address vulnerabilities (Neuhaus et al., 2007b; Shin et al., 2011; Yamaguchi et al., 2012b; Grieco et al., 2016; Wu et al., 2017; Li, Zou, Xu, Ou, Jin, Wang, Deng & Zhong, 2018; Li, Zou, Xu, Jin, Zhu & Chen, 2018; Chernis & Verma, 2018; Huang & Tan, 2019; Li et al., 2019; Lin et al., 2020; Zagane et al., 2020; Li, Zou, Xu, Jin, Zhu & Chen, 2021; Huang & Yu, 2021; Li, Wang & Nguyen, 2021; Huang et al., 2021; Hin et al., 2022; Huang & White, 2022; Fu & Tantithamthavorn, 2022; Aumpansub & Huang, 2022; Huang, 2022; Cao et al., 2022; Huang et al., 2022). Some of them learn characteristics of vulnerabilities using various human-defined features such as source code text features in the source code  (Chernis & Verma, 2018), complexity, code churn, and developer activity metrics (Shin et al., 2011), abstract syntax trees (Yamaguchi et al., 2012b), function imports and function calls (Neuhaus et al., 2007b). The main drawback of these techniques is that they requires considerable human effort to define these features.
Recent techniques use deep learning on the code of programs to detect vulnerabilities so that no human experts is needed to define features (Li, Zou, Xu, Ou, Jin, Wang, Deng & Zhong, 2018; Li, Zou, Xu, Jin, Zhu & Chen, 2018; Li et al., 2019; Zagane et al., 2020; Li, Zou, Xu, Jin, Zhu & Chen, 2021; Hin et al., 2022; Li, Wang & Nguyen, 2021; Fu & Tantithamthavorn, 2022; Aumpansub & Huang, 2022; Cao et al., 2022). Most of them either rely on one type of training data or use imbalanced training data. Our work differs from them by using a balanced dataset combined from different types of training data.
Rather than using static information collected from program code, some techniques apply machine learning on dynamic information collected from the sequences of function calls to detect vulnerabilities (Grieco et al., 2016; Wu et al., 2017). Particularly deep learning models have been shown to have better accuracy than traditional machine learning models (Wu et al., 2017).

### v3

@@ 1 Introduction
As manually finding vulnerabilities typically incur considerable effort and time, numerous studies have been dedicated to automatically identify vulnerabilities (Kim et al., 2017; Li et al., 2016; Grieco et al., 2016a; Neuhaus et al., 2007a; Yamaguchi et al., 2013, 2012a). Primarily, these approaches rely on code similarity detection or pattern matching techniques. However, code similarity detection may not effectively identify vulnerabilities that do not result from code duplication, and pattern matching necessitates the expertise of human professionals to define vulnerability patterns.
@@ 5 Related Work
Many approaches have been proposed to detect and address vulnerabilities (Valeur et al., 2005; Neuhaus et al., 2007b; Dessiatnikoff et al., 2011; Shin et al., 2011; Yamaguchi et al., 2012b; Zheng & Zhang, 2013; Huang & Lie, 2014; Grieco et al., 2016a; Li et al., 2016; Wu et al., 2017; Huang & Lie, 2017; David et al., 2018; Li, Zou, Xu, Ou, Jin, Wang, Deng & Zhong, 2018; Li, Zou, Xu, Jin, Zhu & Chen, 2018; Chernis & Verma, 2018; Wang et al., 2010; Huang & Tan, 2019; Li et al., 2019; Lin et al., 2020; Zagane et al., 2020; Li, Zou, Xu, Jin, Zhu & Chen, 2021; Huang & Yu, 2021; Li, Wang & Nguyen, 2021; Huang et al., 2021; Eshghie et al., 2021; Hin et al., 2022; Huang & White, 2022; Fu & Tantithamthavorn, 2022; Aumpansub & Huang, 2022; Cao et al., 2022). They can be broadly categorized as rule-based approaches and learning-based approaches.
Rule-based approaches detect the existence of vulnerabilities using predefined rules, which typically characterize vulnerable and non-vulnerable program code structures (David et al., 2018) or behaviors (Wang et al., 2010). Rule-based approaches follow the predefined rules to analyze program code or program behaviors. These analyses can be performed dynamically (Huang & Yu, 2021), which execute target programs, or statically (Zheng & Zhang, 2013), which examine target programs without executing them. Rule-based approaches identify a vulnerability when a predefined rule finds a match of vulnerable code structures or behaviors. A major disadvantage of rule-based approaches is that the predefined rules require considerable manual effort and time to generate.
As learning-based approaches have thrived in a myriad of areas, particularly in software security and reliability (Mickens et al., 2007; Yuan et al., 2011; Huang & Lie, 2014; Grieco et al., 2016b; Wang et al., 2016; Long & Rinard, 2016; Huang & Lie, 2017; Cummins et al., 2018; Li et al., 2019; Tien et al., 2020), they have also been leveraged in vulnerability detection. Learning-based approaches extract characteristics of program code or behaviors automatically and identify vulnerabilities based on these characteristics. Conventional machine learning approaches learn characteristics of vulnerabilities using various human-defined features such as source code text features in the source code  (Chernis & Verma, 2018), complexity, code churn, and developer activity metrics (Shin et al., 2011), abstract syntax trees (Yamaguchi et al., 2012b), function imports and function calls (Neuhaus et al., 2007b). Similar to rule-based approaches, a main drawback of conventional machine learning approaches is that they require considerable human effort to define these features.
Recent approaches use deep learning on program code to detect vulnerabilities so that no human experts is needed to define features (Li, Zou, Xu, Ou, Jin, Wang, Deng & Zhong, 2018; Li, Zou, Xu, Jin, Zhu & Chen, 2018; Li et al., 2019; Zagane et al., 2020; Li, Zou, Xu, Jin, Zhu & Chen, 2021; Hin et al., 2022; Li, Wang & Nguyen, 2021; Fu & Tantithamthavorn, 2022; Aumpansub & Huang, 2022; Cao et al., 2022). They typically use neural networks to automatically build classification models from a large number of program samples. Deep learning approaches have been shown to have better accuracy than conventional machine learning approaches (Wu et al., 2017). However, most of them either rely on one type of training data or use imbalanced training data. Our work differs from them by using a balanced dataset that combines different types of training data.